Suppose the value of the S&P 500 Stock Index is currently $3,250.

a. If the one-year T-bill rate is 5.5% and the expected dividend yield on the S&P 500 is 5.0%, what should the one-year maturity futures price be? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

b. What would the one-year maturity futures price be, if the T-bill rate is less than the dividend yield, for example, 4.0%? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

Futures price=Spot*e^((risk free rate-dividend yield)*t)

1.

=3250*e^((5.5%-5%)*1)=3416.63

2.

=3250*e^((4%-5%)*1)=3217.66